solve the expression for x; (4-2)(x-2)=4x-8? can you help.thank you .
Answer by checkley71(8403)

(4-2)(x-2)=4x-8 or 2(x-2)=4x-8 or 2x-4=4x-8 or 2x-4x=4-8 or -2x=-4 or x=2
proof (4-2)(2-2)=4(2)-8 or 2(0)=8-8 or 0=0